# A Multimodal Manufacturing Safety Chatbot: Knowledge Base Design, Benchmark Development, and Evaluation of Multiple RAG Approaches

### v1

@@ 4 Artifact Design and Development
In the third stage of our DSR methodology, we designed and developed a concrete artifact. While many digital solutions could theoretically address the challenge of delivering accessible, accurate, and adaptive safety training in manufacturing, few can simultaneously meet the requirements of accuracy, low latency, and affordability identified in Section 3. For example, generic e-learning platforms or static digital manuals often fall short because they lack contextual awareness, interactivity, and real-time responsiveness. In contrast, modern AI technologies, such as a multimodal chatbot capable of processing text, images, and speech, offer a uniquely flexible and human-centered medium for safety communication. In particular, by embedding RAG techniques within a multimodal interface, such a system can ground its responses in authoritative safety documents while allowing workers to query information naturally through the mode most convenient to their environment. This convergence of adaptive retrieval, multimodal reasoning, and conversational accessibility positions a RAG-based chatbot as a suitable artifact to operationalize the design requirements and realize the goals of Industry 5.0–aligned, human-centric safety training. We next describe the first two phases of our development framework (Figure 3), which focus on data curation and the construction of RAG methods.
@@ 4.1 Phase 1: Source Curation and Corpus Development
Phase 1 establishes the foundation of the chatbot development framework by focusing on the systematic curation and preparation of safety-critical source materials. This stage ensures that all subsequent modeling and retrieval processes are grounded in authoritative, high-quality data. It involves collecting and organizing diverse safety documents into a unified corpus suitable for RAG. This phase includes four steps: (1) collecting the source documents, (2) extracting and cleaning text, (3) segmenting the content into meaningful units, and (4) standardizing how each segment is saved.
@@ 4.1.1 Collect Safety Documents
Grounding LLM responses in authoritative documents reduces hallucinations and improves factual correctness when compared to generation from the LLM alone [lewis2020rag]. We curated four categories of documents to ensure that the chatbot’s responses are grounded in authoritative safety expectations: (1) OSHA Laws and Regulations, i.e., CFR 1910 Subparts O and N, (2) OSHA technical guidance documents, (3) NIOSH safety alerts, and (4) OEM manuals for our target machines. Figure 2 summarizes our data sources.
Phase 1: Source Curation and Corpus Development 8 OSHA Regulations (29 CFR 1910):
@@
@  OSHA Technical Documents:
@@
. NIOSH Safety Alerts and Guidance:
@@ 4.1.2 Extract and Clean Text
For each document, we first determined whether it was machine-readable (i.e., searchable or selectable in a standard PDF reader). For machine-readable PDFs, we used the PyMuPDF library in Python [artifex2025pymupdf], which returns page-level text while preserving paragraph structure. If the document stored text as embedded images, we applied the Mistral OCR API [mistral2025ocr] to produce a text layer with page boundaries and paragraph breaks. After OCR, these documents were processed identically to machine-readable files, ensuring a consistent representation across the corpus.
@@ 4.1.4 Filename Standardization
Once segmented, each section was saved as its own PDF, and any repeated header, footer, or boilerplate content was removed when present. We standardized filenames to include the source type (e.g., OSHA, NIOSH, OEM), the document identifier, the section title, and the original page range. This naming convention ensures that each file is traceable back to its authoritative source and simplifies manual review, debugging, and alignment in downstream retrieval steps.
@@ 4.2 Phase 2: Knowledge Base and Indexing
The goal of this phase is to construct the knowledge bases from the segmented documents in Phase 1 and develop the retrieval mechanisms that the chatbot relies on when answering questions. Prior research has shown that different retrieval strategies can surface different types of relevant content [lewis2020rag]. Because manufacturing safety questions may use precise terminology or broader hazard descriptions, we examined multiple retrieval methods rather than relying on a single strategy. In later phases, we evaluate which approach performs best when answering the benchmark questions developed in Phase 4.
@@ 4.2.1 Cloud-Based Keyword and Semantic Search Indices
During ingestion, we specified a chunk size of 4,000 tokens (approximately 3,000 words; the default is 800 tokens) to ensure that the API did not further subdivide our segmented documents. The OpenAI Retrieval API supports two retrieval modes over the same stored data openai2025retrieval. In keyword search mode, the system prioritizes exact or near-exact lexical matches between the query terms and the stored content. In semantic search mode, the system rewrites or embeds the query and retrieves passages whose vector representations are close in meaning, even when they do not share surface-level wording. Keyword search is most effective when users know and reuse technical terms verbatim, whereas semantic search better supports users who describe their information needs in everyday language or with partial recall of the original phrasing.
@@ 4.2.2 Local Index for Keyword Search
In addition to the cloud-based indices, we constructed a local keyword-based index using the BM25 retriever in the Python LangChain library [langchain2025home]. BM25 employs probabilistic term weighting to rank documents based on estimated relevance and has remained a competitive baseline in information retrieval [robertson2009probabilistic]. We selected BM25 for three primary reasons. First, as a purely lexical method, BM25 is well aligned with queries that contain distinctive regulatory language, component identifiers, or procedural step names, all of which tend to appear verbatim in technical manuals. Second, maintaining the BM25 index locally eliminates per-query API costs and reduces dependency on external services, which is advantageous for repeated experimentation and for potential deployment in cost-sensitive or bandwidth-constrained environments. Third, local execution provides consistently low latency and full control over the indexing and retrieval pipeline, facilitating reproducibility and fine-grained inspection of retrieval behavior.
@@ 4.2.3 Graph-Based Indices
Graph-based retrieval methods allow the system to use both textual similarity and structural relationships among safety concepts, procedures, and component interactions. This is especially valuable in technical domains where relevant information spans multiple steps or subcomponents, and where safe outcomes depend on understanding dependencies rather than isolated passages. Following the LangChain GraphRetriever design [langchain2025graph], we implemented three graph RAG variants that differ in how they explore or constrain graph neighborhoods. AstraDB serves as the online vector storage backend for all three configurations [ibm2025astradb] because it provides scalable embedding storage with native support for vector similarity search and integrates smoothly with graph-based retrieval workflows.
The three variants provide complementary strategies. The Graph Eager retriever expands outward through connected nodes to gather broader contextual information, which helps when queries are imprecise. The Graph MMR variant applies maximal marginal relevance to balance relevance and diversity, reducing redundancy across similar segments. Finally, the vanilla similarity search baseline over AstraDB performs traditional $k$-nearest-neighbor retrieval without graph traversal and serves as a reference for assessing the added value of graph-based methods.
@@ 5 Artifact Demonstration
The fourth stage of our DSR methodology is to “demonstrate the use of the artifact to solve one or more instances of the problem[peffers2007design]. For our case, this involved deploying a publicly accessible multimodal chatbot interface that exposes the safety-assistance capabilities of the underlying RAG system to end users. The system focuses on three types of equipment: the Bridgeport Manual Mill, TL-1 CNC Machine, and the Universal Robots Collaborative Robot (Cobot). The interface supports three input modes: (1) text-based dialogue, (2) image-based queries, and (3) speech-to-speech interaction.
Operationally, the fourth stage corresponds to Phase 3 of our development framework. We began by developing a text chat interface that allows workers to pose safety-related questions in everyday language. The system then retrieves and synthesizes answers grounded in the authoritative documents curated earlier in the pipeline (see Section 4.1). We next added an image input mode, enabling users to upload photographs of equipment, controls, or workspace conditions. The model interprets the visual context and issues a retrieval query tailored to the situation, returning relevant safety passages and procedures. Finally, we added a speech modality to support hands-free interaction, allowing users to speak their questions and receive spoken responses.
The deployed system was implemented as a Streamlit application [streamlit2025v] that integrates OpenAI’s Responses and Realtime APIs to provide text, image, and speech capabilities through a unified interface. Retrieval operates over the same corpus, vector stores, and indices established in earlier phases, preserving traceability between responses and their documentary sources. The application also exposes the retrieval router described in Section 4.2.4, allowing users and researchers to select among different RAG configurations and to experiment with them in situ. This capability supports both practitioner use and systematic comparison in later evaluation phases.
The system logs session-level analytics (e.g., query types, retrieval configurations, and response characteristics) and provides transparent citations linking each answer to specific source passages, supporting explainability, auditability, and user trust. A screenshot of the deployed application is shown in Figure 3, and the full code implementation is available in the project repository (see “Data and Code Availability” section).
@@ 6 Artifact Evaluation
The fifth stage of our DSR methodology comprises “comparing the objectives of a solution to actual observed results from use of the artifact in the demonstration[peffers2007design]. In our context, this stage corresponds to a systematic assessment of how well the proposed RAG configurations meet the design requirements identified earlier, namely high accuracy, low latency, and cost. To enable this assessment, we adopt a three-stage evaluation pipeline within our development framework. In Phase 4, we curate a domain-specific benchmark of safety-related queries and ground-truth reference answers derived from the authoritative corpus. In Phase 5, we use the above benchmark to conduct controlled experiments over multiple RAG configurations. Finally, in Phase 6, we conduct expert-based subjective evaluations to evaluate the most promising configurations, ultimately narrowing down to a single configuration that best balances performance, cost, and latency for deployment.
@@ 6.1 Phase 4: Benchmark Q&A Dataset Development
To evaluate the performance of the RAG systems, we developed a comprehensive benchmark consisting of questions and gold-standard answers derived from the operational and safety manuals of the three target machines. This benchmark was designed to assess the chatbot’s ability to retrieve, reason about, and accurately convey safety-critical information across diverse levels of difficulty.
@@
The process began with our research team conducting a detailed review of each machine’s official manuals. From these materials, we systematically extracted and formulated questions that represent realistic scenarios faced by operators and trainees in manufacturing settings. Each question was paired with a corresponding answer grounded explicitly in the manuals or regulatory guidelines, ensuring factual correctness and traceability. For instance, an example of an easy question might be: What is a pinch hazard in robot operation? with the corresponding answer: A pinch hazard occurs when body parts can be caught between moving robot parts or surfaces. Easy questions can typically be answered from a single section of a manual. To better simulate realistic industrial information retrieval tasks, we also created more challenging questions that required integrating knowledge from multiple sections or documents. For example, a complex question might ask: What are the external connection ports on the robot? The answer should be Teach Pendant Port, SD Card slot, Ethernet, USB 2.0 and USB 3.0, Mini Displayport, 10A Mini Blade Fuse, which requires synthesizing details scattered across different technical sections. This ensured the benchmark tested both direct retrieval and reasoning capabilities of the system.
@@
After the initial Q&A pairs were derived, we implemented a validation phase to ensure content accuracy and industrial relevance. Each question and answer was reviewed either by experienced operators and/or safety professionals from partner companies as well as by researchers with expertise in manufacturing, engineering technology, occupational safety and human factors engineering. Their feedback was used to refine the clarity, technical precision, and contextual relevance of the benchmark items. This iterative validation process guaranteed that the benchmark accurately represented the complexity of real-world safety knowledge and could be reliably used for evaluating AI system performance across different machines and hazard contexts. We refer to the final, validated answers in our benchmarks as gold-standard answers.
@@
In total, we created 60 questions for the UR5 Cobot, 51 for the Haas TL-1 lathe, and 42 for the Bridgeport Manual Mill. Due to space constraints, we henceforth focus on the UR5 Cobot in our evaluations since it exhibits greater operational complexity and a higher degree of human–machine interaction than the other two machines. We argue that the UR5 benchmark sufficiently tests the system’s ability to retrieve and synthesize information across multiple text segments (“chunks”) produced during corpus pre-processing. As such, including additional results for other machines would provide little to no insight beyond what is demonstrated through the Cobot evaluation.
@@ 6.2 Phase 5: Automated Evaluation and Experimental Design
We used a full-factorial experimental design to study how retrieval method and model specification choices influenced the quality of answers generated by the AI. The purpose was to identify which configuration patterns consistently yielded correct responses obtained fast and cheaply before advancing a small number of candidate configurations for human evaluation in Phase 6. In particular, we vary three factors, summarized in Table 1, resulting in $24$ retrieval-generation pipelines. Henceforth, we use the term pipeline to refer to a specific combination of the three factor levels used to generate an answer from the LLM/chatbot. Each pipeline is evaluated across all 60 benchmark questions, yielding $24\times 60=1,440$ generated responses per evaluation metric.
@@
Max Output Tokens Constrains verbosity and potential reasoning detail for generated responses $\{5000\}$, a fixed factor to simplify our experiment
Reasoning Effort Sets structured reasoning depth for LLM $\{low\}$, a fixed factor to simplify our experiment
@@
For each pipeline, we generate an answer to every benchmark question using that pipeline’s specific settings. All outputs are stored together with their pipeline, question, and replicate identifiers to support consistent evaluation. We used a single system prompt (set of instructions) in our experiments. We set the underlying LLM’s temperature parameter to 0 to reduce the randomness from the sampling process, making responses more deterministic for a given pipeline.
Given the generated answers from each pipeline alongside the gold-standard answers, we computed a set of evaluation metrics corresponding to the three design requirements in Section 3. The complete set of metrics and their definitions are summarized in Table 2. First, accuracy was assessed using an LLM-as-judge that compares each generated answer to the gold-standard answer and returns a binary decision. The LLM-based judgment used standardized evaluation prompt templates provided in the LangChain “RAG Evaluation” tutorial [langchain2025evaluate] and ChatGPT 5 as the model evaluator. Second, latency was measured as the end-to-end response time for answer generation. Third, cost values were built into the experiment design by restricting model selection to smaller/distilled versions of ChatGPT 5, namely gpt-5-mini-2025-08-07 and gpt-5-nano-2025-08-07. These LLMs are considered to be high-performing and relatively low-cost. As of October 2025, the gpt-5-mini-2025-08-07 was priced at \$0.25/1M input tokens and \$2/1M output tokens, and gpt-5-nano-2025-08-07 was priced at \$0.050/1M input tokens and \$0.400/1M output tokens [openai2025pricing].
@@
Cost Token usage cost for gpt-5-mini and gpt-5-nano

### v2

@@ 4 Artifact Design and Development
In the third stage of our DSR methodology, we designed and developed a concrete artifact. While many digital solutions could theoretically address the challenge of delivering accessible, accurate, and adaptive safety training in manufacturing, few can simultaneously meet the requirements of accuracy, low latency, and affordability identified in Section 3. For example, generic e-learning platforms or static digital manuals often fall short because they lack contextual awareness, interactivity, and real-time responsiveness. In contrast, modern AI technologies, such as a multimodal chatbot capable of processing text, images, and speech, offer a uniquely flexible and human-centered medium for safety communication. In particular, by embedding RAG techniques within a multimodal interface, such a system can ground its responses in authoritative safety documents while allowing workers to query information naturally through the mode most convenient to their environment. Given the advantages of conversational accessibility, multimodal capability, and ability to ground responses in authoritative documents, a RAG-based chatbot is a suitable artifact to operationalize the design requirements and realize the goals of Industry 5.0-aligned, human-centric safety training. We next describe the first two phases of our framework (Figure 3), which focus on data curation and the construction of RAG methods.
@@ 4.1 Phase 1: Source Curation and Corpus Development
Phase 1 establishes the foundation of the chatbot development framework by focusing on the systematic curation and preparation of safety-critical source materials. This stage ensures that all subsequent modeling and retrieval processes are grounded in authoritative, high-quality data.
@@ 4.1.1 Collect Safety Documents
Grounding LLM responses in authoritative documents reduces hallucinations and improves factual correctness when compared to generation from the LLM alone [Lewis et al., 2020]. We curated four categories of documents to ensure that the chatbot’s responses are grounded in authoritative safety expectations: (1) OSHA Laws and Regulations, i.e., CFR 1910 Subparts O and N, (2) OSHA technical guidance documents, (3) NIOSH safety alerts, and (4) OEM manuals for our target machines. Figure 2 summarizes our data sources.
Phase 1: Source Curation and Corpus Development OSHA Regulations (29 CFR 1910):
@@
OSHA Technical Documents:
@@
NIOSH Safety Alerts and Guidance:
@@ 4.1.2 Extract and Clean Text
For each document, we first determined whether it was machine-readable. For machine-readable PDFs, we used the Python library PyMuPDF, which returns page-level text while preserving paragraph structure. If the document stored text as embedded images, we applied the Mistral OCR API to produce a text layer with page boundaries and paragraph breaks. After OCR, these documents were processed identically to machine-readable files to ensure a unified format across the corpus.
@@ 4.1.4 Filename Standardization
Once segmented, each section was saved as its own PDF, and any repeated header, footer, or boilerplate content was removed when present. We standardized filenames to include the source type, the document identifier, the section title, and the original page range. This naming convention ensures that each file is traceable back to its authoritative source and simplifies manual review, debugging, and alignment in downstream steps.
@@ 4.2 Phase 2: Knowledge Base and Indexing
The goal of this phase is to construct the knowledge bases from the segmented documents in Phase 1 and develop the retrieval mechanisms that the chatbot relies on when answering questions. Prior research has shown that different retrieval strategies can surface different types of relevant content [Lewis et al., 2020]. Because manufacturing safety questions may use precise terminology or broader hazard descriptions, we examined multiple retrieval methods rather than relying on a single strategy. In later phases, we evaluate which approach performs best when answering the benchmark questions developed in Phase 4.
@@ 4.2.1 Cloud-Based Keyword and Semantic Search Indices
During ingestion, we specified a chunk size of 4,000 tokens (approximately 3,000 words; the default is 800 tokens) to ensure that the API did not further subdivide our segmented documents. As of Feb. 06, 2026, the OpenAI Retrieval API supports two retrieval modes over the same stored data. In keyword search mode, the system prioritizes exact or near-exact lexical matches between the query terms and the stored content. In semantic search mode, the system rewrites or embeds the query and retrieves passages whose vector representations are close in meaning, even when they do not share surface-level wording. Keyword search is most effective when users know and reuse technical terms verbatim, whereas semantic search better supports users who describe their information needs in everyday language or with partial recall of the original phrasing.
@@ 4.2.2 Local Index for Keyword Search
We constructed a local keyword-based index using the BM25 retriever in the Python LangChain library. BM25 employs probabilistic term weighting to rank documents based on estimated relevance and has remained a competitive baseline in information retrieval [Robertson and Zaragoza, 2009]. We selected BM25 for three primary reasons. First, as a purely lexical method, BM25 is well aligned with queries that contain distinctive regulatory language, component identifiers, or procedural step names, all of which tend to appear verbatim in technical manuals. Second, maintaining the BM25 index locally eliminates per-query API costs and reduces dependency on external services, which is advantageous for repeated experimentation and for potential deployment in cost-sensitive or bandwidth-constrained environments. Third, local execution provides consistently low latency and full control over the indexing and retrieval pipeline, facilitating reproducibility and fine-grained inspection of retrieval behavior.
@@ 4.2.3 Graph-Based Indices
Graph-based retrieval methods allow the system to use both textual similarity and structural relationships among safety concepts, procedures, and component interactions. This is especially valuable in technical domains where relevant information spans multiple steps or subcomponents, and where safe outcomes depend on understanding dependencies rather than isolated passages. Following the LangChain GraphRetriever design [LangChain, 2025b], we implemented three graph RAG variants that differ in how they explore or constrain graph neighborhoods. AstraDB serves as the online vector storage backend for all three configurations because it provides scalable embedding storage with native support for vector similarity search and integrates smoothly with graph-based retrieval workflows.
The three variants provide complementary strategies. The Graph Eager retriever expands outward through connected nodes to gather broader context, which helps when queries are imprecise. The Graph MMR variant applies maximal marginal relevance to balance relevance and diversity, reducing redundancy across similar segments. Finally, the vanilla similarity search baseline over AstraDB performs traditional $k$-nearest-neighbor retrieval without graph traversal and serves as a reference for assessing the added value of graph methods.
@@ 5 Artifact Demonstration
The fourth stage of our DSR methodology is to “demonstrate the use of the artifact to solve one or more instances of the problem[Peffers et al., 2007]. For our case, this involved deploying a publicly accessible multimodal chatbot interface that exposes the safety-assistance capabilities of the underlying RAG system to end users. The system focuses on three types of equipment: the Bridgeport Manual Mill, TL-1 CNC Machine, and the Universal Robots Collaborative Robot (Cobot). The interface supports three input modes: (1) text-based dialogue, (2) image-based queries, and (3) speech-to-speech interaction.
Operationally, the fourth stage corresponds to Phase 3 of our development framework. We began by developing a text chat interface that allows workers to pose safety-related questions in plain language. The system then retrieves and synthesizes answers grounded in the authoritative documents curated earlier in the pipeline (see Section 4.1). We next added an image input mode, enabling users to upload photographs of equipment, controls, or workspace conditions. The model interprets the visual context and issues a retrieval query tailored to the situation, returning relevant safety passages. Finally, we added a speech modality to enable hands-free questions and spoken responses.
The deployed system was implemented as a Streamlit application that integrates OpenAI’s Responses and Realtime APIs to provide text, image, and speech capabilities through a unified interface. Retrieval operates over the same corpus, vector stores, and indices established in earlier phases, preserving traceability between responses and their documentary sources. The application also exposes the retrieval router described in Section 4.2.4, allowing users and researchers to select among different RAG configurations and to experiment with them in situ. This capability supports both practitioner use and systematic comparison in later evaluation phases.
The system logs session-level analytics and provides transparent citations linking each answer to specific source passages, supporting explainability, auditability, and user trust. A screenshot of the deployed application is shown in Figure 3, and the full code implementation is available in the project repository (see “Data and Code Availability” section).
@@ 6 Artifact Evaluation
The fifth stage of our DSR methodology comprises “comparing the objectives of a solution to actual observed results from use of the artifact in the demonstration[Peffers et al., 2007]. In our context, this stage corresponds to a systematic assessment of how well the proposed RAG configurations meet the design requirements identified earlier, namely high accuracy, low latency, and cost. To enable this assessment, we adopt a three-stage evaluation pipeline within our development framework. In Phase 4, we curate a domain-specific benchmark of safety-related queries and ground-truth reference answers derived from the authoritative corpus. In Phase 5, we use the above benchmark to conduct controlled experiments over multiple RAG configurations. Finally, in Phase 6, we conduct expert-based subjective evaluations to evaluate the most promising configurations, ultimately narrowing down to a single configuration that best balances performance, cost, and latency for deployment.
@@ 6.1 Phase 4: Benchmark Q&A Dataset Development
We developed a comprehensive benchmark consisting of questions and gold-standard answers derived from the operational and safety manuals of the three target machines. This benchmark was designed to assess the chatbot’s ability to retrieve, reason about, and accurately convey safety-critical information across diverse levels of difficulty.
@@
The process began with our research team conducting a detailed review of each machine’s official manuals. From these materials, we systematically extracted and formulated questions that represent realistic scenarios faced by operators and trainees in manufacturing settings. Each question was paired with a corresponding answer grounded explicitly in the manuals or regulatory guidelines, ensuring factual correctness and traceability. For instance, an example of an easy question was: What is a pinch hazard in robot operation? with the corresponding answer: A pinch hazard occurs when body parts can be caught between moving robot parts or surfaces. Easy questions can typically be answered from a single section of a manual. To better simulate realistic industrial information retrieval tasks, we also created more challenging questions that required integrating knowledge from multiple sections or documents. For example, a complex question might ask: What are the external connection ports on the robot? The answer should be Teach Pendant Port, SD Card slot, Ethernet, USB 2.0 and USB 3.0, Mini Displayport, 10A Mini Blade Fuse, which requires synthesizing details scattered across different technical sections. This ensured the benchmark tested direct retrieval and reasoning capabilities of the system.
@@
After the initial Q&A pairs were derived, we implemented a validation phase to ensure content accuracy and industrial relevance. Each Q&A was reviewed either by experienced safety professionals from partner companies as well as by researchers with expertise in manufacturing, engineering technology, occupational safety and human factors engineering. Their feedback was used to refine the clarity, technical precision, and contextual relevance of the benchmark items. This iterative validation process guaranteed that the benchmark accurately represented the complexity of real-world safety knowledge and could be reliably used for evaluating AI system performance across different machines and hazard contexts. We refer to the final, validated answers in our benchmarks as gold-standard answers.
@@
In total, we created 60 questions for the UR5 Cobot, 51 for the Haas TL-1 lathe, and 42 for the Bridgeport Manual Mill. Due to space constraints, we henceforth focus on the UR5 Cobot in our evaluations since it exhibits greater operational complexity and a higher degree of human–machine interaction than the other two machines. We argue that the UR5 benchmark sufficiently tests the system’s ability to retrieve and synthesize information across multiple text segments (“chunks”) produced during corpus pre-processing. As such, including additional results for other machines would provide little to no insight beyond what is demonstrated through the Cobot evaluation. Our UR5 dataset contains 23, 30, and 7 questions classified as easy, moderate, and challenging, respectively.
@@ 6.2 Phase 5: Automated Evaluation and Experimental Design
We used a full-factorial experimental design to study how retrieval method and model specification choices influenced the quality of answers generated by the AI. The purpose was to identify which configuration patterns consistently yielded correct responses obtained fast and cheaply before advancing a small number of candidate configurations for human evaluation in Phase 6. We vary three factors (see Table 1), resulting in $24$ retrieval-generation pipelines. Henceforth, we use the term pipeline to refer to a specific combination of the three factor levels used to generate an answer from the LLM/chatbot. Each pipeline is evaluated across all 60 benchmark questions, yielding $24\times 60=1,440$ generated responses per evaluation metric.
@@
For each pipeline, we generate an answer to every benchmark question using that pipeline’s specific settings. All outputs are stored together with their pipeline, question, and replicate identifiers to support consistent evaluation. We used a single system prompt (set of instructions) in our experiments. The prompt explicitly instructs the model to use only the retrieved sources when generating answers, preventing it from relying on prior knowledge. To promote concise and comparable outputs, the prompt instructs the model to provide direct answers in one or two sentences when possible and to add brief supporting sentences only when additional explanation is necessary. We set the underlying LLM’s temperature parameter to 0 to ensure reproducibility, making responses deterministic for a given pipeline. Moreover, we set the maximum number of output tokens to $5,000$ to avoid truncation of safety-critical answers while still keeping costs bounded, and we set the reasoning effort to lowlow to avoid introducing additional latency and cost.
Given the generated answers from each pipeline alongside the gold-standard answers, we computed a set of evaluation metrics corresponding to the three design requirements in Section 3. The complete set of metrics and their definitions are summarized in Table 2. First, accuracy was assessed using an LLM-as-judge that compares each generated answer to the gold-standard answer and returns a binary decision. The LLM-based judgment used standardized evaluation prompt templates provided in the LangChain “RAG Evaluation” tutorial [LangChain, 2025a] and ChatGPT 5 as the model evaluator. Second, latency was measured as the end-to-end response time for answer generation. Third, cost values were built into the experiment design by restricting model selection to smaller/distilled versions of ChatGPT 5, namely gpt-5-mini-2025-08-07 and gpt-5-nano-2025-08-07. These LLMs are considered to be high-performing and relatively low-cost. As of October 2025, the gpt-5-mini-2025-08-07 was priced at \$0.25/1M input tokens and \$2/1M output tokens, and gpt-5-nano-2025-08-07 was priced at \$0.050/1M input tokens and \$0.400/1M output tokens.
@@
Cost Token cost for gpt-5-mini and gpt-5-nano